Five years ago, Grace was 2 less than twice the age of jones. 27 years from now, Grace will be 2 more than thrice the age of Jones next year. What are their present ages?

Answer by stanbon(75887)

Five years ago, Grace was 2 less than twice the age of jones. 27 years from now, Grace will be 2 more than thrice the age of Jones next year. What are their present ages?
-----
Equations:
g-5 = 2(j-5)-2
g+27 = 3(j+1)+2
------------------------
Rearrange::
g - 2j = -7
g - 3j = -22
----------------
Subtract and solve for "j"::
j = 15 yrs (Jones's age now)
g = 2j-7
g = 23 yrs (Grace's age now)
--------------
Cheers,
Stan H.
--------------